# A Certifiable Algorithm for Simultaneous Shape Estimation and Object Tracking

### v2

@@ I Introduction
footnotetext: This work was partially supported by the AFOSR “Certifiable and Self-Supervised Category-Level Tracking” program, Carlone’s NSF CAREER award, the ONR RAPID program, and by MathWorks.
@@
A significant body of work is dedicated to object pose and shape estimation in the single-frame case [5, 6, 7, 8]. However, using a single image for detection and estimation sacrifices temporal information which is readily available. A good tracking algorithm leverages temporal consistency to operate in noisy, occlusion-rich, or highly dynamic environments. When the object shape is known exactly, instance-level tracking algorithms leverage known geometry [9, 4] or extensive training on the specific object instance [3, 10, 11]. However, in practical applications the object is rarely known exactly. Autonomous vehicles must track any of the thousands of car models in their surroundings; even if all were to be cataloged, the vehicle must still reason over intra-category variations.
@@
Recent work has made considerable progress towards category-level and category-free object tracking using end-to-end learning [12, 13] or learning-based segmentation combined with local solvers for estimation [14, 15, 16]. While often effective in practice, learning-based approaches have unpredictable failure modes when used outside their training domain, while local-solver-based approaches are prone to converge to local minima corresponding to poor estimates. For safety critical applications, it is important to have interpretable and predictable models that produce provably optimal estimates.
Contribution. We propose the first certifiably optimal algorithm for category-level shape estimation and pose tracking (Fig. 1). In our problem formulation (Section III), we consider as input the 3D position of keypoints detected in an RGB-D image sequence. These inputs are fed to a fixed-lag smoother that performs estimation of the states over a receding horizon, while enforcing smoothness of the motion using a constant-twist motion model. We parametrize the object’s shape using an active shape model [17, 18], which describes shape as a linear combination of 3D models from a library of representative models. Our first key contribution (Section IV) is to show that despite the non-convexity of the estimation problem, we can develop an empirically tight semidefinite relaxation that computes certifiably optimal object poses, velocities, and shape without the need for an initial guess. We name the resulting approach CAST (Certifiable Algorithm for Shape estimation and Tracking). Our second contribution (Section V) is to extend CAST to the practical case where some of the keypoint measurements are outliers; we handle this case by combining an outlier-pruning method with a robust estimator based on graduated non-convexity [19]. We call the resulting outlier-robust approach CAST#. Our final contribution (Section VI) is to evaluate our methods in extensive experiments, and show they produce accurate estimates when tested in simulation, on the YCBInEOAT dataset [3], and in a drone-based vehicle tracking scenario.
@@ II Related Work
Category-Level and Category-Free Object Tracking. In practical settings, instance-level information is rarely available. Recent approaches investigate pose and shape estimation for objects within a known category [14, 25, 12, 16, 13] or at least similar enough to the training data [15, 26]. These approaches generally extract a sparse representation of the object to estimate relative motion between frames. Wang et al. [14] focus on an attention mechanism for extracting frame-to-frame keypoints in a self-supervised manner, leaving the work of relative pose estimation to point cloud registration, which is unable to use temporal information beyond two frames. Wen and Bekris [25] use a similar architecture but take a SLAM-inspired approach, using dense frame-to-frame feature correspondences and multi-frame pose graph optimization to refine the estimate. Other methods use learned keypoint correspondences for the Iterative Closest Point (ICP) method [13] or learning-based regression to estimate relative motion in the small pose regime [12]. Even with keyframe selection [15], frame-to-frame back-ends require a separate tool to obtain object pose relative to a camera or world frame, which is often useful in applications. In contrast, we propose an optimization back-end that produces certifiably optimal shape and pose estimates from category-level keypoints without relying on local solvers. This gives useful world-frame poses directly and allows the use of a motion model to mitigate the impact of measurement noise.
Certifiable Algorithms. Our work extends the body of work on certifiable perception algorithms: a certifiable algorithm solves an optimization problem and either provides a certificate of optimality or a bound on the suboptimality of the produced solution [9]. Certifiable algorithms are typically derived using semidefinite relaxations, and are usually based on Shor’s relaxation of Quadratically Constrained Quadratic Programs (QCQP) or Lasserre’s relaxation of polynomial optimization problems [27, 28, 29]. Certifiable algorithms have been proposed for rotation averaging [30, 31], pose graph optimization [32, 33, 34], 3D registration [35, 9], 2-view geometry [36, 37, 38], perspective-n-point problems [39], and single-frame pose and shape estimation [5]. Recent work has extended certifiable solvers to cope with outliers [27] and anisotropic noise [40]. Our approach extends [5] to tracking over a receding time horizon using a motion model.
@@ III Problem Formulation
In the following, we assume that the velocity noise follows an isotropic zero-mean Gaussian distribution: ${\bm{v}}_{t}^{\epsilon}\sim\mathcal{N}(0,{\bm{\Sigma}}_{t}^{v})$. Moreover, we assume that the relative rotation noise follows an isotropic Langevin distribution about the identity, following standard practice [32] for distributions over $\mathrm{SO}(3)$: $\mathbf{R}^{\epsilon}_{t}\sim\mathcal{L}(\mathbf{I}_{3},\kappa_{t})$. In this equation, $\kappa_{t}$ is the concentration parameter of the Langevin distribution (intuitively, this plays a similar role as the inverse of the variance).
@@
We use the active shape model to represent intra-category shape variations. Given an object category (e.g.,bottle), we assume a library of 3D models (e.g. specific bottle shapes) that span the category, where the objects in the library are denoted as $\mathcal{B}_{k}$, $k=1,\ldots,K$. Any instance, then, is just a pointwise linear combination of the models in the shape library (see Fig. 2). More formally, let $\mathbf{x}_{i}$ be any point on the instance object that corresponds to the point $\mathbf{b}_{i}^{k}\in\mathcal{B}_{k}$ in each library shape, $k=1,\ldots,K$. The active shape model is:
@@
where $c_{k}\in[0,1]$ and $\sum_{k}c_{k}=1$. Thus, the shape of the target object is fully specified by its shape coefficient $\mathbf{c}=[c_{1},\ldots,c_{K}]$ and the shape library for each point $\mathbf{B}_{i}=[\mathbf{b}_{i}^{1},\ldots,\mathbf{b}_{i}^{K}]$. This representation is simple and expressive: it captures any object in the convex hull of the shape library (including the library shapes themselves) via a linear combination described by a single vector of coefficients [17, 18]. Further, measurements of a small number of semantic keypoints are enough to resolve the object shape.
@@
Problem 1 (Shape Estimation and Pose Tracking)
@@ IV CAST: Certifiable Shape Estimation and Tracking in the Outlier-Free Setting
In the previous expession, we used the shorthand $\Delta\mathbf{c}\triangleq\mathbf{c}-\mathbf{\bar{c}}$ and assumed isotropic covariances ${\bm{\Sigma}}_{t}^{i}\triangleq\frac{1}{w_{t}^{i}}{\mathbf{I}}_{3}$ and ${\bm{\Sigma}}_{t}^{v}\triangleq\frac{1}{w_{t}}{\mathbf{I}}_{3}$. We also relaxed the constraint $c_{k}\geq 0$. We observe that the objective is the sum of the shape prior with the negative log-likelihoods of the measurements (4) and dynamics (2). The constraints enforce the domains of the variables (e.g. $\mathbf{R}_{t}\in\mathrm{SO}(3)$ or $\mathbf{1}_{K}^{\mathsf{T}}\mathbf{c}=1$) and the dynamics (1). We prove that (5) is a maximum a posteriori estimator in Appendix A.
@@ IV-A Closed-Form Solution for Shape
Proposition 1 (Optimal Shape)
For any choice of positions and rotations $(\mathbf{p}_{t},\mathbf{R}_{t})$, the optimal shape coefficient solving (5) is
@@
Proposition 1 is proven in Appendix B. This insight is similar to the one used in [5] for the case of static object.
@@ IV-B Change of Variables to Quadratic Program
First, we observe that the constraint set is already quadratic: the dynamics (1) are quadratic equalities and $\mathrm{SO}(3)$ constraints on rotations can be written as quadratic equality constraints, see, e.g., [42]. Moreover, the expression of $\mathbf{c^{\star}}$ we found in (6) is also quadratic. Some of the terms in the objective, however, are not quadratic. Particularly, $\mathbf{R}_{t}\mathbf{B}_{i}\mathbf{c^{\star}}$ is quadratic and appears in a squared norm, leading to a quartic expression. Fortunately, the rotational invariance of the $\ell_{2}$ norm allows a simple reparameterization of the position variable as $\mathbf{s}_{t}\triangleq\mathbf{R}^{\mathsf{T}}_{t}\mathbf{p}_{t}$, turning the problem into a QCQP. This result is summarized below.
Proposition 2 (QCQP Formulation)
The shape estimation and tracking problem can be equivalently written as a quadratically constrained quadratic program (QCQP):
@@
where $\mathbf{c}$ is defined as in (6) and is a linear function of $\mathbf{R}_{t}$ and $\mathbf{s}_{t}$.
The proof of the proposition is given in Appendix C.
@@ IV-C Convex Semidefinite Relaxation
Corollary 3 (Shor’s Semidefinite Relaxation)
@@
Similar to relaxations derived in related work [32, 46, 5] the rank of $\mathbf{X^{\star}}$ is a certificate for the optimality of the solution. Moreover, we can bound the suboptimality of a feasible solution to (10) obtained from (11) using the objective. Given a feasible solution $(\hat{\mathbf{x}},\hat{\mathbf{v}})$ achieving objective $\hat{f}$ in (10), we bound its suboptimality using $\hat{f}-f^{\star}_{\mathrm{SDP}}\geq\hat{f}-f^{\star}\geq 0$. The condition $\hat{f}=f^{\star}_{\mathrm{SDP}}$ also certifies the optimality of the solution. The scalar $d=f^{\star}-f^{\star}_{SDP}$ is typically referred to as the duality gap, while the difference $\hat{f}-f^{\star}_{\textrm{SDP}}$ is sometimes called the relaxation gap.
The relaxation (11) is relevant in practice because we observe it to be empirically tight in the case of low-to-moderate noise and no outliers; hence it can produce optimal solutions without needing an initial guess. Moreover, the size of the SDP (11) is independent of the size of the shape library, hence the relaxation is relatively efficient to solve in practice. We name the resulting approach CAST: Certifiable Algorithm for Shape estimation and Tracking.
@@ V-A Compatibility Checks to Remove Gross Outliers
Inspired by [5], we introduce compatibility tests to identify gross outliers. These tests rely on the assumptions of rigid-body motion of the object and the active shape model. The most likely inlier set is thus the largest set of compatible measurements, found via a fast mixed-integer linear program.
@@
Proposition 4 (Shape Compatibility Test)
@@
Proposition 5 (Time Compatibility Test)
@@
Proposition 6 (Largest Set of Compatible Measurements)
@@
$\mathop{\arg\!\min}\limits_{{\bm{\theta}}\in\{0,1\}^{N\times T}}$ $\sum_{t=1}^{T}\sum_{i=1}^{N}\theta_{t}^{i}$ (15)
@@ VI Experiments
This section characterizes CAST and CAST# in synthetic and real settings. Synthetic experiments (Section VI-A) show that the semidefinite relaxation in CAST is empirically tight and returns accurate estimates in the presence of measurement and process noise, while CAST# is robust to 50-60% outliers. Section VI-B shows that CAST#is competitive with other category-level approaches on YCBInEOAT [3], a manipulation focused dataset for object tracking. Lastly, we show that CAST# remains accurate in a real-world drone-based vehicle tracking scenario (Section VI-C).
@@ VI-A Optimality and Robustness in Synthetic Dataset
Dataset. We generate keypoint measurements according to the measurements and motion models in Section III. Ground truth trajectories follow the constant twist model (2) with Gaussian velocity noise and Langevin rotation rate noise (process noise). The ground truth trajectory and randomly generated shape determine the measured keypoint positions without regard for occlusion, subject to Gaussian perturbations (measurement noise) and outliers. We use the realistic PASCAL3D+ aeroplane shape library [48] (with characteristic length $l=0.2$ m) to generate a ground truth shape vector. In each experiment, we fix measurement noise to $5$% of the characteristic length, and the process noise to $0.01$m and $0.01$rad. For the measurement noise experiment, we set the velocity weights $\omega_{t}=1$ to improve tightness; we provide results with standard weights in Appendix E.
@@
Baselines. We compare CAST against PACE [5], a certifiably optimal solver for single-frame pose estimation, and PACE+EKF, an approach that filters the pose estimate from PACE at each time using an extended Kalman filter (EKF) while using a constant-twist motion model [49]. We test CAST with a time horizon of $4$, $8$, or $12$ frames and label the corresponding results as CAST-4, CAST-8, and CAST-12; we also report CAST-U, which is a variant of CAST-12 with no velocity or rotation smoothing ($\omega_{t}=0$, $\kappa_{t}$ = 0). Finally, to test CAST#, we replace a fraction of the measurements with random outliers normally distributed about the centroid of the object with a standard deviation equal to the characteristic length. For the tests with outliers, we compare against ablations of CAST# with only GNC or only compatibility-based (MILP) outlier rejection without GNC and use $T=12$. In each experiment we show the median and interquartile range of the error of the last estimated state over 500 runs.
@@
Results. Fig 4 reports the median position error (as a percent of length scale), rotation error (in degrees), shape error, and suboptimality gap or solve time for increasing measurement noise (normalized by length scale), process noise (reported as a multiple of $5$%), and outlier ratio. In all experiments, CAST and CAST# achieve the lowest median position, rotation, and shape error. In Figs. 4(a) and 4(b), CAST is consistently tight (suboptimality gap $<10^{-4}$) in low to moderate noise, and still gives an accurate estimate when not tight. Interestingly, while CAST outperforms its unsmoothed variant CAST-U, the latter remains tight for higher measurement noise. CAST-U benefits over PACE from access to additional measurements, and CAST benefits over CAST-U from additional information about the object’s motion. The primary cost of CAST compared to PACE is its runtime. In our tests, the runtime ranged from 0.1 to 7 Hz depending on the time horizon (Appendix E). We also note the poor performance of PACE-EKF in both experiments. The EKF provides some benefit for very low noise but quickly diverges for higher noise as the distribution of PACE measurements deviates from Gaussian and the dynamics are nonlinear. We provide a comparison to an EKF using perturbed ground truth poses for the update step in Appendix E.

### v3

@@ Abstract
This paper has been accepted for publication in IEEE Robotics and Automation Letters. Please cite the paper as: Lorenzo Shaikewitz, Samuel Ubellacker, and Luca Carlone, “A Certifiable Algorithm for Simultaneous Shape Estimation and Object Tracking” IEEE Robotics and Automation Letters, Dec. 2024.
@@ I Introduction
A significant body of work is dedicated to object pose and shape estimation in the single-frame case [5, 6, 7]. However, using a single image for detection and estimation sacrifices temporal information which is readily available. A good tracking algorithm leverages temporal consistency to operate in noisy, occlusion-rich, or highly dynamic environments. When the object shape is known exactly, instance-level tracking algorithms leverage known geometry [8, 4] or extensive training on the specific object instance [3, 9, 10]. However, in practical applications the object is rarely known exactly. Autonomous vehicles must track any of the thousands of car models in their surroundings; even if all were to be cataloged, the vehicle must still reason over intra-category variations.
@@
Recent work has made considerable progress towards category-level and category-free object tracking using end-to-end learning [11, 12] or learning-based segmentation combined with local solvers for estimation [13, 14, 15]. While often effective in practice, learning-based approaches have unpredictable failure modes when used outside their training domain, while local-solver-based approaches are prone to converge to local minima corresponding to poor estimates. Safety-critical applications demand interpretable and predictable models that produce provably optimal estimates.
Contribution. We propose the first certifiably optimal algorithm for category-level shape estimation and pose tracking (Fig. 1). In our problem formulation (Section III), we consider as input the 3D position of keypoints detected in an RGB-D image sequence. These inputs are fed to a fixed-lag smoother that performs estimation of the states over a receding horizon, while enforcing smoothness of the motion using a constant-twist motion model. We parametrize the object’s shape using an active shape model [16, 17], which describes shape as a linear combination of 3D models from a library of representative models. Our first key contribution (Section IV) is to show that despite the non-convexity of the estimation problem, we can develop an empirically tight semidefinite relaxation that computes certifiably optimal object poses, velocities, and shape without the need for an initial guess. We name the resulting approach CAST (Certifiable Algorithm for Shape estimation and Tracking). Our second contribution (Section V) is to extend CAST to the practical case where some of the keypoint measurements are outliers; we handle this case by combining an outlier-pruning method with a robust estimator and graduated non-convexity [18]. We call the resulting outlier-robust approach CAST#. Our final contribution (Section VI) is to evaluate our methods in extensive experiments, and show they produce accurate estimates when tested in simulation, on the YCBInEOAT [3] and NOCS [7] datasets, and in a drone-based vehicle tracking scenario.
@@ II Related Work
Category-Level and Category-Free Object Tracking. In practical settings, instance-level information is rarely available. Recent approaches investigate pose and shape estimation for objects within a known category [13, 24, 11, 15, 12] or at least similar enough to the training data [14]. These approaches generally extract a sparse representation of the object to estimate relative motion between frames. Wang et al. [13] focus on an attention mechanism for extracting frame-to-frame keypoints in a self-supervised manner, leaving the work of relative pose estimation to point cloud registration, which is unable to use temporal information beyond two frames. Wen and Bekris [24] use a similar architecture but take a SLAM-inspired approach, using dense frame-to-frame feature correspondences and multi-frame pose graph optimization to refine the estimate. Other methods use learned keypoint correspondences for the Iterative Closest Point (ICP) method [12] or learning-based regression to estimate relative motion in the small pose regime [11]. Even with keyframe selection [14], frame-to-frame back-ends require a separate tool to obtain object pose relative to a camera or world frame, which is often useful in applications. In contrast, we propose an optimization back-end that produces certifiably optimal shape and pose estimates from category-level keypoints without relying on local solvers. This gives useful world-frame poses directly and allows the use of a motion model to mitigate the impact of measurement noise.
Certifiable Algorithms. Our work extends the body of work on certifiable perception algorithms: a certifiable algorithm solves an optimization problem and either provides a certificate of optimality or a bound on the suboptimality of the produced solution [8]. Certifiable algorithms are typically derived using semidefinite relaxations, and are usually based on Shor’s relaxation of Quadratically Constrained Quadratic Programs (QCQP) or Lasserre’s relaxation of polynomial optimization problems [25, 26, 27]. Certifiable algorithms have been proposed for rotation averaging [28, 29], pose graph optimization [30, 31], 3D registration [32, 8], 2-view geometry [33, 34], perspective-n-point problems [35], and single-frame pose and shape estimation [6]. Recent work has extended certifiable solvers to cope with outliers [25] and anisotropic noise [36]. Our approach extends [6] to tracking over a receding time horizon using a motion model.
@@ III Problem Formulation
In the following, we assume that the velocity noise follows an isotropic zero-mean Gaussian distribution: ${\bm{v}}_{t}^{\epsilon}\sim\mathcal{N}(0,{\bm{\Sigma}}_{t}^{v})$ and that the relative rotation noise follows an isotropic Langevin distribution about the identity, following standard practice [30] for distributions over $\mathrm{SO}(3)$: $\mathbf{R}^{\epsilon}_{t}\sim\mathcal{L}(\mathbf{I}_{3},\kappa_{t})$. In this equation, $\kappa_{t}$ is the concentration parameter of the Langevin distribution (intuitively, this plays a similar role as the inverse of the variance).
@@
We use the active shape model to represent intra-category shape variations. Given an object category (e.g.,  bottle), we assume a library of 3D models (e.g.,  specific bottle shapes) that span the category, where the objects in the library are denoted as $\mathcal{B}_{k}$, $k=1,\ldots,K$. Any instance, then, is just a pointwise linear combination of the models in the shape library (see Fig. 2). More formally, let $\mathbf{x}_{i}$ be a point on the instance object corresponding to the point $\mathbf{b}_{i}^{k}\in\mathcal{B}_{k}$ in each library shape. The active shape model is:
@@
where $c_{k}\in[0,1]$ and $\sum_{k}c_{k}=1$. Thus, the shape of the target object is fully specified by its shape coefficient $\mathbf{c}=[c_{1},\ldots,c_{K}]$ and the shape library for each point $\mathbf{B}_{i}=[\mathbf{b}_{i}^{1},\ldots,\mathbf{b}_{i}^{K}]$. This representation is simple and expressive: it captures any object in the convex hull of the shape library (including the library shapes themselves) via a linear combination described by a single vector of coefficients [16, 17]. Further, measurements of a small number of semantic keypoints are enough to resolve the dense object shape.
@@
Problem 1.
@@ IV CAST: Certifiable Shape Estimation and Tracking in the Outlier-Free Setting
In the previous expession, we used the shorthand $\Delta\mathbf{c}\triangleq\mathbf{c}-\mathbf{\bar{c}}$ and assumed isotropic covariances ${\bm{\Sigma}}_{t}^{i}\triangleq\frac{1}{w_{t}^{i}}{\mathbf{I}}_{3}$ and ${\bm{\Sigma}}_{t}^{v}\triangleq\frac{1}{w_{t}}{\mathbf{I}}_{3}$. We also relaxed the constraint $c_{k}\geq 0$. We observe that the objective is the sum of the shape prior with the negative log-likelihoods of the measurements (4) and dynamics (2). The constraints enforce the domains of the variables (e.g. $\mathbf{R}_{t}\in\mathrm{SO}(3)$ or $\mathbf{1}_{K}^{\mathsf{T}}\mathbf{c}=1$) and the dynamics (1). Eq. (5) is a maximum a posteriori estimator; see Appendix A for proof.
@@ IV-A Closed-Form Solution for Shape
Proposition 1 (Optimal Shape).
For any positions and rotations $(\mathbf{p}_{t},\mathbf{R}_{t})$, the optimal shape coefficient solving (5) is
@@
Proof.
Similar to [6], see Appendix B. ∎
@@ IV-B Change of Variables to Quadratic Program
First, we observe that the constraint set is already quadratic: the dynamics (1) are quadratic equalities and $\mathrm{SO}(3)$ constraints on rotations can be written as quadratic equality constraints, see, e.g., [38]. In the objective, however, the squared norm of $\mathbf{R}_{t}\mathbf{B}_{i}\mathbf{c^{\star}}$ is quartic. We use the rotational invariance of the $\ell_{2}$ norm to reparametrize position as $\mathbf{s}_{t}\triangleq\mathbf{R}^{\mathsf{T}}_{t}\mathbf{p}_{t}$, turning the problem into a QCQP. The result is summarized below.
Proposition 2 (QCQP Formulation).
Let $\mathbf{c}$ be defined as in (6), and note that it is a linear function of $\mathbf{R}_{t}$ and $\mathbf{s}_{t}$. The shape estimation and tracking problem can be equivalently written as a quadratically constrained quadratic program:
@@
Proof.
See Appendix C. ∎
@@ IV-C Convex Semidefinite Relaxation
Corollary 3 (Shor’s Relaxation).
@@
Similar to relaxations derived in related work [30, 42, 6] the rank of $\mathbf{X^{\star}}$ is a certificate for the optimality of the solution. Moreover, we can bound the suboptimality of a feasible solution to (10) obtained from (11) using the objective. Given a feasible solution $(\hat{\mathbf{x}},\hat{\mathbf{v}})$ achieving objective $\hat{f}$ in (10), we bound its suboptimality using $\hat{f}-f^{\star}_{\mathrm{SDP}}\geq\hat{f}-f^{\star}\geq 0$. The condition $\hat{f}=f^{\star}_{\mathrm{SDP}}$ also certifies the optimality of the solution. The scalar $\hat{f}-f^{\star}_{\textrm{SDP}}$ is called the suboptimality gap.
The relaxation (11) is relevant in practice because we observe it to be empirically tight in the case of low-to-moderate noise and no outliers; hence it can produce optimal solutions without needing an initial guess. Moreover, the size of the SDP (11) is independent of the size of the shape library, hence the relaxation is relatively efficient to solve. We name the resulting approach CAST: Certifiable Algorithm for Shape estimation and Tracking.
@@ V-A Compatibility Checks to Remove Gross Outliers
Inspired by [6], we introduce compatibility tests to identify gross outliers (Fig. 3). These tests rely on the assumptions of rigid-body motion of the object and the active shape model. The most likely inlier set is thus the largest set of compatible measurements, found via a fast mixed-integer linear program.
@@
Proposition 4 (Shape Compatibility Test).
@@
Proposition 5 (Time Compatibility Test).
@@
Proposition 6 (Largest Set of Compatible Measurements).
@@
$\mathop{\arg\!\max}\limits_{{\bm{\theta}}\in\{0,1\}^{N\times T}}$ $\sum_{t=1}^{T}\sum_{i=1}^{N}\theta_{t}^{i}$ (15)
@@ VI Experiments
This section characterizes CAST. Synthetic experiments (Section VI-A) show the semidefinite relaxation in CAST is empirically tight and returns accurate estimates in the presence of noise, while CAST# is robust to 50-60% outliers. Sections VI-B, VI-C, and VI-D show CAST#is competitive with other category-level approaches on two public datasets and a real-world drone-based vehicle tracking scenario.
@@ VI-A Optimality and Robustness in Synthetic Dataset
Dataset. We generate keypoint measurements according to the measurements and motion models in Section III. Ground truth trajectories follow the constant twist model (2) with Gaussian velocity noise and Langevin rotation rate noise (process noise). The ground truth trajectory and randomly generated shape determine the measured keypoint positions without regard for occlusion, subject to Gaussian perturbations (measurement noise) and outliers. We use the realistic PASCAL3D+ aeroplane shape library [44] (with characteristic length $l=0.2$ m) to generate a ground truth shape vector. In each experiment, we fix measurement noise to $5$% of the characteristic length, and the process noise to $0.01$ m and $0.01$ rad. For the measurement noise experiment, we set the velocity weights $\omega_{t}=1$ to improve tightness; results with standard weights are provided in Appendix E.
@@
Baselines. We compare CAST against PACE [6], a certifiably optimal solver for single-frame pose estimation, and PACE+EKF, an approach that filters the pose estimate from PACE at each time using an extended Kalman filter (EKF) while using a constant-twist motion model. We test CAST with a time horizon of $4$, $8$, or $12$ frames and label the corresponding results as CAST-4, CAST-8, and CAST-12; we also report CAST-U, which is a variant of CAST-12 with no velocity or rotation smoothing ($\omega_{t}=0$, $\kappa_{t}$ = 0). Finally, to test CAST#, we replace a fraction of the measurements with random outliers normally distributed about the centroid of the object with a standard deviation equal to the characteristic length. For the tests with outliers, we compare against ablations of CAST# with only GNC or only compatibility-based (MILP) outlier rejection without GNC and use $T=12$.
@@
Results. Fig 4 reports the median position error (as a percent of length scale), rotation error (in degrees), shape error ($l_{2}$ distance between predicted and ground truth shape vector $\mathbf{c}$), and suboptimality gap or solve time for increasing measurement noise (normalized by length scale), process noise (reported as a multiple of $5$%), and outlier ratio. In all experiments, CAST and CAST# achieve the lowest median position, rotation, and shape error. In Figs. 4(a) and 4(b), CAST is consistently tight (suboptimality gap $<10^{-4}$) in low to moderate noise, and still gives an accurate estimate when not tight. Interestingly, while CAST outperforms its unsmoothed variant CAST-U, the latter remains tight for higher measurement noise. CAST-U benefits over PACE from access to additional measurements, and CAST benefits over CAST-U from additional information about the object’s motion. The primary cost of CAST compared to PACE is its runtime, which ranged from 0.1 to 7 Hz depending on the time horizon; see Appendix E for detailed runtimes.
We also note the poor performance of PACE-EKF in both experiments. The EKF provides some benefit for very low noise but quickly diverges for higher noise as the distribution of PACE measurements deviates from Gaussian and the dynamics are nonlinear. We provide a comparison to an EKF using perturbed ground truth poses in Appendix E.